# One-cycle Structured Pruning with Stability Driven Structure Search

### v1

@@ 4.2 VGG16/19 on CIFAR-10/100
Tab. 1 compares the pruning of VGG16/19 on CIFAR-10/100 datasets using our method against several state-of-the-art approaches. The VGG16 pruning on CIFAR-10 is compared with RCP [34], LAASP [18], PGMFP [4], CPGCN [26], OTO [5, 6], DLRFC [24], and DCFF [39]. Among these pruning methods, like our method, OTO and DCFF also pruned the network from scratch. While OTO, CPGCN, and DLRFC achieve higher parameter reduction rates, their top-1 accuracy is lower than ours. Notably, OTO uses a specially designed training optimizer, whereas our method works with the standard SGD optimizer. With only $26.01\%$ and $21.22\%$ of the original network FLOPs remaining, we achieved top-1 accuracy of $93.88\%$ and $93.76\%$, respectively—the highest reported for similar pruning rates. Notably, while methods like DLRFC and DCFF show improvements in pruned model performance, this is largely due to their comparison against relatively lower baseline accuracy.
Tab. 1 also shows our method’s VGG19 pruning performance in comparison with EigenDamage [54], GReg [56], and DepGraph [13] on the CIFAR-100 dataset. We achieved the highest top-1 accuracy of $70.47\%$ while reducing the network parameters around $95\%$ and FLOPs around $89\%$. Although GReg [56] also utilizes the concept of growing regularization and uses a fully trained network for pruning, its top-1 accuracy is nearly 3% lower than ours.
@@
VGG16/ CIFAR-10 RCP [34] 47.99 42.11 94.33 93.94 0.39
PGMPF [4] 34.00 - 93.68 93.60 0.08
CPGCN [26] 26.93 6.94 93.10 93.08 0.02
LAASP [18] $\times$ 39.54 26.83 93.79 93.79 0.00
OTOv1 [5] $\times$ 26.80 5.50 93.20 93.30 -0.10
$\text{OCSPruner{}{}}_{74\%}$$\times$ 26.01 9.72 94.08 93.88 0.20
[…]
OTOv2 [6] $\times$ 23.70 4.90 93.20 93.20 0.00
DCFF [39] $\times$ 23.13 7.20 93.02 93.47 -0.45
$\text{OCSPruner{}{}}_{79\%}$$\times$ 21.22 7.46 94.08 93.76 0.32
VGG19/ CIFAR-100 ED [54] 11.37 - 73.35 65.18 8.17
GReg-2 [56] 11.31 - 74.02 67.75 6.27
DepGraph [13] 11.21 - 73.50 70.39 3.11
$\text{OCSPruner{}{}}_{89\%}$$\times$ 11.23 5.49 74.24 70.47 3.77
@@ 4.3 ResNet on CIFAR-10
The pruning of ResNet on CIFAR-10 using our method consistently outperforms state-of-the-art methods across depths 56 and 110 (Tab. 2). While our proposed algorithm automatically learns slimmer sub-networks from scratch, we not only compare its performance with other automatic methods like FPGM [23], and LAASP [18] but also assess its effectiveness against various other state-of-the-art techniques such as DLRFC [24], HRank [36], ResRep [12], DCFF [39], DepGraph [13], GReg [56], C-SGD [10], and Rethink [43]. Similar to ours, LAASP [18] employs a partially trained network for pruning. However, it ignores structurally coupled parameters while estimating filter saliency scores and adapts manually defined stable pruning epochs. In contrast to our technique, which gradually moves pruning parameters towards zero, FPGM [23] directly sets filters with low saliency scores to zero, while still allowing updates in the next training epoch. We prune ResNet56 and ResNet110 models for two different FLOPs reduction rates. Notably, with $38.88\%$ reduced FLOPs, ResNet56 achieves $93.65\%$ top-1 accuracy. Again, ResNet110 achieves a remarkable $94.13\%$ top-1 accuracy with only 33.10% FLOPs.
@@
Depth Method
Pre-
train?
@@
56 HRank [36] 50.00 - 93.26 93.17 0.09
DLRFC [25] 47.42 44.37 93.06 93.57 -0.51
ResRep [12] 47.09 - 93.71 93.71 0.00
NuSPM [29] $\times$ 49.72 - - 93.50 -
FPGM [23] $\times$ 47.40 - 93.59 92.93 0.66
LAASP [18] $\times$ 47.40 61.43 93.61 93.45 0.16
$\text{OCSPruner{}{}}_{53\%}$$\times$ 46.93 50.27 93.97 93.80 0.17
[…]
C-SGD [10] 39.15 - 93.39 93.44 -0.05
DepGraph [13] 38.91 - 93.53 93.64 -0.11
DCFF [39] $\times$ 43.75 44.71 93.26 93.26 0.00
OTOv2 [6] $\times$ 38.88 - -
$\text{OCSPruner{}{}}_{61\%}$$\times$ 38.88 41.42 93.97 93.65 0.32
110 FPGM [23] $\times$ 47.70 - 93.68 93.74 -0.06
$\text{OCSPruner{}{}}_{54\%}$$\times$ 46.37 52.71 94.36 94.30 0.06
[…]
LAASP [18] $\times$ 41.29 54.56 94.41 93.61 0.80
DCFF [39] $\times$ 33.18 32.37 93.50 93.80 -0.30
$\text{OCSPruner{}{}}_{67\%}$$\times$ 33.10 34.93 94.36 94.13 0.23
@@
Pre-
train?
FLOPs
Params
@@
ResNet50 CLR-RNF [37] 59.60 66.20 76.01 74.85 1.16 92.96 92.31 0.65 0.66$\leq$ 0.66$\times$ 0.66 ×
FPGM [23] $\times$ 57.80 - 76.15 75.03 1.12 92.87 92.40 0.47 1.00$\leq$ 1.00$\times$ 1.00 ×
LAASP [18] $\times$ 57.67 77.29 76.48 75.85 0.63 93.14 92.81 0.33 1.04×1.04$\times$1.04 ×
[…]
Tayler [45] 54.95 - 76.18 74.50 1.68 - - - <0.50×<0.50$\times$< 0.50 ×
White-Box [64] 54.34 68.63 76.15 75.32 0.83 92.96 92.43 0.53 0.66$\leq$ 0.66$\times$ 0.66 ×
DepGraph [13] 51.82 - 76.15 75.83 0.32 - - - <0.50×<0.50$\times$< 0.50 ×
[…]
PGMPF [4] 46.50 - 76.15 75.11 0.90 92.93 92.41 0.52 0.66$\leq$ 0.66$\times$ 0.66 ×
GReg-2 [56] 43.33 62.30 76.13 75.36 0.77 92.86 92.41 0.45 0.66$\leq$ 0.66$\times$ 0.66 ×
PaT [52] $\times$ 41.30 - - 74.85 - - - - >1.00×>1.00$\times$> 1.00 ×
NuSPM [29] $\times$ 41.30 - - 75.30 - - - - >1.00×>1.00$\times$> 1.00 ×
[…]
DTP [35] 32.68 - 76.13 74.26 1.87 - - - 0.66×{0.$6$
OTOv1 [5] $\times$ 34.50 35.50 76.10 74.70 1.40 92.90 92.10 0.80 <1.00×<1.00$\times$< 1.00 ×
[…]
MobileNetV2 NetAda [59] 72.73 - 72.00 70.00 2.00 - - - -
CC [33] 71.67 - 71.88 70.91 0.97 - - - 0.66$\leq$ 0.66$\times$ 0.66 ×
AMC [22] $\times$ 70.00 - 71.80 70.80 1.00 - - - -
[…]
DCFF [39] $\times$ 46.67 74.86 72.00 68.60 3.40 90.12 88.13 1.99 >1.00×>1.00$\times$> 1.00 ×
LAASP [18] $\times$ 45.50 64.09 71.79 68.45 3.34 90.47 88.40 2.07 1.08×1.08$\times$1.08 ×
[…]
@@ 4.4 ResNet50/MobileNetV2 on ImageNet
We evaluate the performance of ResNet50 and MobileNetV2 on the ImageNet dataset across different pruning ratios, comparing our results with those of various state-of-the-art methods from the literature (Tab. 3). Our method consistently surpasses the performance of several state-of-the-art approaches. Notably, similar to our approach, FPGM [23], PaT [52], NuSPM [29], and OTOv1 [5] also achieve a fully trained pruned network using one-cycle training from scratch. Remarkably, with a mere $43\%$ of the original FLOPs remaining in the network, our algorithm achieves the highest accuracy ($75.49\%$ top-1 and $92.63\%$ top-5) among state-of-the-art methods in the literature for similar pruning ratios.
Fig. 1 graphically illustrates the relationship between network FLOPs and top-1 accuracy, providing an approximation of the training speed-up associated with each pruning method in reference to the baseline network training. This clearly shows that our proposed technique consistently achieves higher top-1 accuracy compared to several other state-of-the-art methods while also being the most efficient pruning framework among them.
Tab. 3 also presents the pruning outcomes for MobileNetV2 on ImageNet and compares them with several state-of-the-art methods, including NetAda [59], CC [33], AMC [22], DCFF [39], LAASP [18], and DepGraph [13]. To ensure a fair comparison, we exclusively consider results for the $1.0$ scale version of MobileNetV2. Experiments are conducted at $30\%$ and $55\%$ pruning ratios to align with commonly used benchmarks. In both instances, our proposed method attains superior top-1 accuracy compared to the other methods.
[…]
@@ 4.5 Ablation Studies
Table 4: The influence on varying Sparsity Learning (SL) start epoch ($t_{\text{sl-start}}$) while pruning ResNet50 on the ImageNet. Given $t_{\text{sl-start}}$, $t^{*}$, i.e., stable pruning epoch, was automatically identified using Eq. 6. The value $t_{\text{sl-start}}=17$ was automatically estimated using Eq. 5, while other values were manually assigned.
@@
For additional results and ablation studies, we direct the readers to the supplementary material of this paper.
5 Conclusion
In this work, we introduce an efficient and stable pruning framework named OCSPruner.Unlike traditional methods that rely on fully trained networks for pruning, our approach integrates pruning early in the training process, achieving competitive performance within a single training cycle. OCSPruner automatically identifies stable pruning architecture in the early training stage guided by the pruning stability indicator. An integral aspect of any pruning procedure lies in determining the optimal sub-network based on parameter saliency criteria. Consequently, our future research will focus on exploring better saliency criteria suitable for our pruning framework, surpassing the current reliance on simple norm-based criteria. Furthermore, we aim to evaluate and adapt our pruning methodology for diverse tasks such as object detection, segmentation, diffusion models, and large language models.
Limitation. While our method effectively employs regularization to encourage the reduction of unimportant network structures, additional attention is required for fine-tuning hyper-parameters tailored to each model architecture and dataset. Moreover, the proposed algorithm is particularly well-suited for learning schedules characterized by gradual variations, such as cosine or linear schedules, enabling a smoother detection of pruning stability by avoiding abrupt changes in learning rates.
Supplementary Material
6 Additional Ablation Studies
6.1 Group Saliency vs Conventional Saliency
Conventionally, filter saliency estimation evaluated each convolutional layer in isolation when ranking filters based on their importance. In contrast, group saliency used in our paper considers all parameters within a group, including those from interconnected convolutional layers, batch normalization layers, and fully connected layers. Also, given the varied dimensions of different group parameters, as mentioned in Sec. 3.2 of the main paper, we performed local normalization before computing group saliency scores. Tab. 5 demonstrates that utilizing group saliency for pruning leads to better performance compared to conventional saliency estimation.
Table 5: Performance comparison of pruned networks on the CIFAR-10 dataset using OCSPruner with conventional versus group saliency.
Model Group Saliency Conventional Saliency
FLOPs (%) Params (%) Acc. (%) FLOPs (%) Params (%) Acc. (%)
Vgg16 19.99 7.18 93.87 20.00 3.75 92.94
ResNet32 50.09 57.88 92.94 50.04 38.61 92.60
ResNet56 46.88 47.66 93.85 47.03 43.93 93.48
6.2 Analysis of Structured Sparsity Regularization
Regularization in the proposed pruning pipeline helps the stability indicator reach a stable epoch for final pruning. In particular, we utilize Eq. 8 and  Eq. 9 from Sec. 3.4 of the main paper for this purpose. While Eq. 8 exhibit a slower sparsity learning, Eq. 9 is designed to speed up this process. Tab. 6 presents accuracy results for comparable FLOPs pruning rates, considering the application of regularization with Eq. 8 and Eq. 9 independently, and in combination. It is important to highlight that the parameter reduction rate is automatically determined based on the pruned architecture for a given FLOPs reduction rate. The accuracy values in the table indicate that better results are achieved when both equations Eq. 8 and Eq. 9 are used together for structured sparsity regularization. Fig. 5 illustrate the progress of sparsity learning, starting from the sparsity learning initiation epoch ($t_{\text{sl-start}}$) up to the epoch of stable pruning ($t^{*}$).
Table 6: Ablation study on pruning accuracy analysis using Eq. 8 and Eq. 9 independently and in combination for structured sparsity regularization.
Model/Dataset
Sparsity
Learning
FLOPs (%)
Params (%)
Top-1 Acc
Top-5 Acc
ResNet32/ CIFAR-10 Eq. 8 44.69 38.59 92.09 -
Eq. 9 44.74 46.94 92.47 -
Eq. 8 and Eq. 9 44.98 48.28 92.74 -
ResNet18/ ImageNet Eq. 8 55.09 55.13 67.89 88.00
Eq. 9 55.08 66.55 68.52 88.31
Eq. 8 and Eq. 9 55.04 67.90 68.66 88.61
6.3 OCSPruner with Pre-trained Models
Although the primary focus of this paper is on pruning models from scratch within a single training cycle, we also applied the proposed pruning framework to pre-trained models, as it is more practical in scenarios where pre-trained models are already available. Under identical training and pruning settings and hyper-parameters in OCSPruner, we replaced the randomly initialized input model with a pre-trained model. The pruning results for VGG16 and ResNet56 on the CIFAR-10 dataset, as well as for the ResNet50 model on the ImageNet dataset, are presented in Tab. 7. The results demonstrate that the algorithm also performs effectively with pre-trained models, showing no significant accuracy differences compared to pruning for randomly initialized models. Further tuning of the hyper-parameters for pruning pre-trained models could potentially enhance performance. This indicates that OCSPruner is effective for both randomly initialized and pre-trained models.
@@ 6.3 OCSPruner with Pre-trained Models
7 Training and Pruning Hyper-parameters

### v2

@@ 4.2 VGG16/19 on CIFAR-10/100
Tab. 1 compares the pruning of VGG16/19 on CIFAR-10/100 datasets using our method against several state-of-the-art approaches. The VGG16 pruning on CIFAR-10 is compared with RCP [li2022revisiting], LAASP [GHIMIRE2023104745], PGMPF [Cai_An_Yang_Yan_Xu_2022], CPGCN [ijcai2022p431], OTO [chen2021only, chen2023otov2], DLRFC [he2022filtereccv], and DCFF [lin2023training]. Among these methods, similar to ours, OTO, and DCFF also pruned the network from scratch. While OTO, CPGCN, and DLRFC achieve higher parameter reduction rates, their top-1 accuracy is lower than ours. Notably, OTO uses a specially designed training optimizer, whereas our method works with the standard SGD optimizer. With only $26.01\%$ and $21.22\%$ of the original network FLOPs remaining, we achieved top-1 accuracy of $93.88\%$ and $93.76\%$, respectively, the highest reported for similar pruning rates. Notably, while DLRFC and DCFF show improvements in pruned model performance, this is largely due to their comparison against relatively lower baseline accuracy.
Tab. 1 also shows our method’s VGG19 pruning performance in comparison with EigenDamage [wang2019eigendamage], GReg [wang2021neural], and DepGraph [fang2023depgraph] on the CIFAR-100 dataset. We achieved the highest top-1 accuracy of $70.47\%$ while reducing the network parameters around $95\%$ and FLOPs around $89\%$. Although GReg [wang2021neural] also utilizes the concept of growing regularization and uses a fully trained network for pruning, its top-1 accuracy is nearly 3% lower than ours.
@@
VGG16/ CIFAR-10 RCP [li2022revisiting] 47.99 42.11 94.33 93.94 0.39
PGMPF [Cai_An_Yang_Yan_Xu_2022] 34.00 - 93.68 93.60 0.08
CPGCN [ijcai2022p431] 26.93 6.94 93.10 93.08 0.02
LAASP [GHIMIRE2023104745] $\times$ 39.54 26.83 93.79 93.79 0.00
OTOv1 [chen2021only] $\times$ 26.80 5.50 93.20 93.30 -0.10
OCSPruner $\times$ 26.01 9.72 94.08 93.88 0.20
DLRFC [he2022filtereccv] 23.05 5.62 93.25 93.64 -0.39
OTOv2 [chen2023otov2] $\times$ 23.70 4.90 93.20 93.20 0.00
DCFF [lin2023training] $\times$ 23.13 7.20 93.02 93.47 -0.45
OCSPruner $\times$ 21.22 7.46 94.08 93.76 0.32
VGG19/ CIFAR-100 ED [wang2019eigendamage] 11.37 - 73.35 65.18 8.17
GReg-2 [wang2021neural] 11.31 - 74.02 67.75 6.27
DepGraph [fang2023depgraph] 11.21 - 73.50 70.39 3.11
OCSPruner $\times$ 11.23 5.49 74.24 70.47 3.77
@@ 4.3 ResNet on CIFAR-10
The pruning of ResNet on CIFAR-10 using our method consistently outperforms state-of-the-art methods across depths 56 and 110 (Tab. 2). While our proposed algorithm automatically learns slimmer sub-networks from scratch, we not only compare its performance with other automatic methods like FPGM [he2019filter], and LAASP [GHIMIRE2023104745] but also assess its effectiveness against various other state-of-the-art techniques such as DLRFC [he2022filtereccv], HRank [lin2020hrank], ResRep [ding2021resrep], ATO [Wu_2024_CVPR], DCFF [lin2023training], DepGraph [fang2023depgraph], GReg [wang2021neural], C-SGD [ding2019centripetal], and Rethink [liu2018rethinking]. Similar to ours, LAASP [GHIMIRE2023104745] employs a partially trained network for pruning. However, it ignores structurally coupled parameters while estimating filter saliency scores and adapts manually defined stable pruning epochs. In contrast to our technique, which gradually moves pruning parameters towards zero, FPGM [he2019filter] directly sets filters with low saliency scores to zero, while still allowing updates in the next training epoch. We prune ResNet56 and ResNet110 models for two different FLOPs reduction rates. Notably, with $38.88\%$ reduced FLOPs, ResNet56 achieves $93.65\%$ top-1 accuracy. Again, ResNet110 achieves a remarkable $94.13\%$ top-1 accuracy with only 33.10% FLOPs.
@@
Depth
Method
Pre- train?
@@
56 HRank [lin2020hrank] 50.00 - 93.26 93.17 0.09
DLRFC [he2022filtereccv] 47.42 44.37 93.06 93.57 -0.51
ResRep [ding2021resrep] 47.09 - 93.71 93.71 0.00
NuSPM [Lee_2024_WACV] $\times$ 49.72 - - 93.50 -
FPGM [he2019filter] $\times$ 47.40 - 93.59 92.93 0.66
LAASP [GHIMIRE2023104745] $\times$ 47.40 61.43 93.61 93.45 0.16
OCSPruner $\times$ 46.93 50.27 93.97 93.80 0.17
GReg-2 [wang2021neural] 39.21 - 93.36 93.36 0.00
C-SGD [ding2019centripetal] 39.15 - 93.39 93.44 -0.05
DepGraph [fang2023depgraph] 38.91 - 93.53 93.64 -0.11
ATO [Wu_2024_CVPR] $\times$ 45.00 - 93.50 93.74 -0.24
DCFF [lin2023training] $\times$ 43.75 44.71 93.26 93.26 0.00
OCSPruner $\times$ 38.88 41.42 93.97 93.65 0.32
110 FPGM [he2019filter] $\times$ 47.70 - 93.68 93.74 -0.06
OCSPruner $\times$ 46.37 52.71 94.36 94.30 0.06
HRank [lin2020hrank] 41.80 - 93.50 93.36 0.14
LAASP [GHIMIRE2023104745] $\times$ 41.29 54.56 94.41 93.61 0.80
DCFF [lin2023training] $\times$ 33.18 32.37 93.50 93.80 -0.30
OCSPruner $\times$ 33.10 34.93 94.36 94.13 0.23
@@
Pre-train?
FLOPs (%)
Params (%)
@@
ResNet50 CLR-RNF [lin2022pruning] 59.60 66.20 76.01 74.85 1.16 92.96 92.31 0.65 $\leq 0.66\times$
FPGM [he2019filter] $\times$ 57.80 - 76.15 75.03 1.12 92.87 92.40 0.47 $\leq 1.00\times$
LAASP [GHIMIRE2023104745] $\times$ 57.67 77.29 76.48 75.85 0.63 93.14 92.81 0.33 $1.04\times$
$\textbf{OCSPruner{}{}}_{43\%}$ $\times$ 56.99 64.45 76.29 76.00 0.29 93.04 92.90 0.14 $1.28\times$}
HRank [lin2020hrank] 56.18 - 76.15 74.98 1.17 92.87 92.33 0.54 $0.50\times$
Tayler [molchanov2019importance] 54.95 - 76.18 74.50 1.68 - - - $0.50\times$
White-Box [zhang2022carrying] 54.34 68.63 76.15 75.32 0.83 92.96 92.43 0.53 $\leq 0.66\times$
DepGraph [fang2023depgraph] 51.82 - 76.15 75.83 0.32 - - - $0.50\times$
$\textbf{OCSPruner{}{}}_{48\%}$ $\times$ 51.74 62.66 76.29 75.95 0.34 93.04 92.75 0.29 $1.30\times$}
GNN-RL [yu2022topology] 47.00 - 76.10 74.28 1.82 - - - $\leq 0.66\times$
PGMPF [Cai_An_Yang_Yan_Xu_2022] 46.50 - 76.15 75.11 0.90 92.93 92.41 0.52 $\leq 0.66\times$
GReg-2 [wang2021neural] 43.33 62.30 76.13 75.36 0.77 92.86 92.41 0.45 $\leq 0.66\times$
PaT [Shen_2022_CVPR] $\times$ 41.30 - - 74.85 - - - - $1.00\times$
NuSPM [Lee_2024_WACV] $\times$ 41.30 - - 75.30 - - - - $1.00\times$
$\textbf{OCSPruner{}{}}_{57\%}$ $\times$ 42.98 51.83 76.29 75.49 0.80 93.04 92.63 0.41 $1.38\times$}
CC [li2021towards] 37.32 41.39 76.15 74.54 1.61 92.87 92.25 0.62 $\leq 0.66\times$
DTP [li2023differentiable] 32.68 - 76.13 74.26 1.87 - - - $0.66\times$}
OTOv1 [chen2021only] $\times$ 34.50 35.50 76.10 74.70 1.40 92.90 92.10 0.80 $1.00\times$
$\textbf{OCSPruner{}{}}_{66\%}$ $\times$ 33.05 37.13 76.29 74.72 1.57 93.04 92.13 0.91 $1.41\times$}
MobileNetV2 CC [li2021towards] 71.67 - 71.88 70.91 0.97 - - - $\leq 0.66\times$
AMC [he2018amc] $\times$ 70.00 - 71.80 70.80 1.00 - - - -
$\textbf{OCSPruner{}{}}_{30\%}$ $\times$ 70.32 87.88 71.79 70.96 0.83 90.47 89.84 0.63 $1.20\times$}
DepGraph [fang2023depgraph] 45.45 - 71.87 68.46 3.41 - - - $0.50\times$
DCFF [lin2023training] $\times$ 46.67 74.86 72.00 68.60 3.40 90.12 88.13 1.99 $1.00\times$
LAASP [GHIMIRE2023104745] $\times$ 45.50 64.09 71.79 68.45 3.34 90.47 88.40 2.07 $1.08\times$
$\textbf{OCSPruner{}{}}_{55\%}$ $\times$ 45.40 67.76 71.79 68.68 3.11 90.47 88.16 2.31 $1.20\times$}
@@ 4.4 ResNet50/MobileNetV2 on ImageNet
We evaluate the performance of ResNet50 and MobileNetV2 on the ImageNet dataset across different pruning ratios, comparing our results with those of various state-of-the-art methods from the literature (Tab. 3). Our method consistently surpasses the performance of several state-of-the-art approaches. Notably, similar to our approach, FPGM [he2019filter], PaT [Shen_2022_CVPR], NuSPM [Lee_2024_WACV], and OTOv1 [chen2021only] also achieve a fully trained pruned network using one-cycle training from scratch. Remarkably, with a mere $43\%$ of the original FLOPs remaining in the network, our algorithm achieves the highest accuracy ($75.49\%$ top-1 and $92.63\%$ top-5) among state-of-the-art methods in the literature for similar pruning ratios. Fig. 1 graphically illustrates the relationship between network FLOPs and top-1 accuracy, providing an approximation of the training speed-up associated with each pruning method in reference to the baseline network training. This clearly shows that our proposed technique consistently achieves higher top-1 accuracy compared to several other state-of-the-art methods while also being the most efficient pruning framework among them.
Tab. 3 also presents the pruning outcomes for MobileNetV2 on ImageNet and compares them with several state-of-the-art methods, including CC [li2021towards], AMC [he2018amc], DCFF [lin2023training], LAASP [GHIMIRE2023104745], and DepGraph [fang2023depgraph]. We report results using the standard $1.0\times$ MobileNetV2 for fair comparison. Experiments are conducted at $30\%$ and $55\%$ pruning ratios to align with commonly used benchmarks. In both instances, our proposed method attains superior top-1 accuracy compared to the other methods.
The last column in Tab. 3 presents the comparison results for training speed enhancement (higher the better) relative to the baseline ResNet50 model training cost. As shown, our proposed technique achieved up to a $1.41\times$} increase in training speed while pruning 67% of ResNet50 FLOPs. We approximated training speedup for state-of-the-art methods based on factors such as the total number of training rounds and whether pruning was done iteratively or in one shot. This comparative speedup analysis demonstrates that our method is the most efficient among the existing techniques.
@@ 4.5 Ablation Studies
Table 4: The influence on varying Sparsity Learning (SL) start epoch ($t_{\text{sl-start}}$) while pruning ResNet50 on the ImageNet. Given $_{\text{sl-start}}, tt^{*}$, i.e., stable pruning epoch, was identified using Eq. 6.
@@
[…]
Automatically estimated using Eq. 5; all other SL Start Epoch values were manually assigned.
Pruning Pre-trained Models. While OCSPruner is primarily designed for pruning from scratch, we also applied it to pre-trained models to reflect practical scenarios. Using the same training and pruning settings, we replaced randomly initialized models with pre-trained ones. As shown in Tab. 5, OCSPruner achieves significant compression on VGG16 and ResNet56 (CIFAR-10) and ResNet50 (ImageNet), with minimal accuracy loss. This shows its effectiveness for randomly initialized and pre-trained models, with potential for further gains via hyperparameter tuning.
@@
5 Conclusion
We presented OCSPruner, a novel one-cycle structured pruning framework that integrates training and pruning into a single efficient process. By introducing a stability-driven sub-network search guided by a pruning stability indicator, our method enables early and effective pruning. Through structured sparsity regularization and adaptive penalty scheduling, OCSPruner ensures both stable pruning and high model accuracy. Experiments across CIFAR and ImageNet benchmarks demonstrate that OCSPruner achieves state-of-the-art performance in terms of accuracy and training efficiency, outperforming several competitive methods. Moreover, it remains versatile and effective for both from-scratch and pre-trained models. Future work will investigate alternative saliency metrics and broader applications to modern architectures and other domains.
Limitation. Although our method removes redundant structures via regularization, it requires careful, task- and architecture-specific hyperparameter tuning. The algorithm is best suited for training schedules with gradual changes (e.g., cosine or linear decay), as they enable stable pruning by avoiding abrupt shifts in learning rates.
Acknowledgment. This work was supported by the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT) (RS-2025-16071992), the convergence security core talent training business support program(IITP-2024 2024-RS-2024-00426853) supervised by the IITP(Institute of Information & Communications Technology Planning & Evaluation), and G-LAMP Program of the National Research Foundation of Korea (NRF) grant funded by the Ministry of Education (No. RS-2025-25441317). Additionally, this work was also supported by the Technology Innovation Program (20018906, Development of autonomous driving collaboration control platform for commercial and task assistance vehicles) funded by the Ministry of Trade, Industry and Energy (MOTIE, Korea).
[…]
Supplementary Material
6 Additional Ablation Studies
6.1 Group Saliency vs Conventional Saliency
Conventionally, filter saliency estimation evaluates each convolutional layer in isolation when ranking filters based on their importance. In this approach, filters are assessed solely based on their individual contributions to the output of their respective layers, without accounting for interdependencies with other components in the network. In contrast, the group saliency used in our paper considers all parameters within a group, including those from interconnected convolutional layers, batch normalization layers, and fully connected layers.
Given the varied dimensions and scales of different group parameters, as described in Sec. 3.2, we apply local normalization before computing the group saliency scores. This step ensures a fair comparison across heterogeneous parameter types and prevents components with inherently larger magnitudes from being overemphasized. As shown in Tab. 6, pruning based on group saliency consistently outperforms conventional saliency estimation, leading to better accuracy-compression trade-offs.
Table 6: Performance comparison on CIFAR-10 using OCSPruner with group vs. conventional saliency.
Model Group Saliency Conventional Saliency
FLOPs
Params
Acc.
FLOPs
Params
Acc.
Vgg16 19.99 7.18 93.87 20.00 3.75 92.94
ResNet32 50.09 57.88 92.94 50.04 38.61 92.60
ResNet56 46.88 47.66 93.85 47.03 43.93 93.48
6.2 Analysis of Structured Sparsity Regularization
Regularization in the proposed pruning pipeline helps the stability indicator reach a stable epoch for final pruning. In particular, we utilize Eq. 8 and  Eq. 9 from Sec. 3.4 of the main paper for this purpose. While Eq. 8 exhibit a slower sparsity learning, Eq. 9 is designed to speed up this process. Tab. 7 presents accuracy results for comparable FLOPs pruning rates, considering the application of regularization with Eq. 8 and Eq. 9 independently, and in combination. It is important to highlight that the parameter reduction rate is automatically determined based on the pruned architecture for a given FLOPs reduction rate. The accuracy values in the table indicate that better results are achieved when both equations Eq. 8 and Eq. 9 are used together for structured sparsity regularization. Fig. 5 illustrates the progress of sparsity learning, starting from the sparsity learning initiation epoch ($t_{\text{sl-start}}$) up to the epoch of stable pruning ($t^{*}$).
Table 7: Ablation study on pruning accuracy using Eq. 8 and Eq. 9 independently and in combination for structured sparsity regularization.
Model/ Dataset Sparsity Learning FLOPs (%) Params (%) Top-1 Acc. (%) Top-5 Acc. (%)
Eq. 8 Eq. 9
ResNet32/ CIFAR-10 $\times$ 44.69 38.59 92.09 -
$\times$ 44.74 46.94 92.47 -
44.98 48.28 92.74 -
ResNet18/ ImageNet $\times$ 55.09 55.13 67.89 88.00
$\times$ 55.08 66.55 68.52 88.31
55.04 67.90 68.66 88.61
7 Training and Pruning Details
7.1 Hyper-parameter Configuration